OpenStudy (anonymous):

Solve this system of linear equation. -9x+2y=-16 19x+3y=41

OpenStudy (anonymous):

it was wrong :/

OpenStudy (anonymous):

i have another question, can you help me? pleaseee :(

OpenStudy (anonymous):

.x=2,y=1

OpenStudy (anonymous):

yea whats the question

OpenStudy (anonymous):

ayynadiax3Star Solve this system of linear equation 4x+9y=-3 8x+12y=12

OpenStudy (anonymous):

multiply the first equation with 2 and subtract the second equation

OpenStudy (anonymous):

so if i multiply it would be, 8x+18y=-6 ?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

what do i do next?

OpenStudy (anonymous):

subtract the second equation from the first equation

OpenStudy (anonymous):

6,-3 ?

OpenStudy (anonymous):

yea u r right

OpenStudy (anonymous):

I understand how to do it now :) thank you <3

OpenStudy (anonymous):

always welcome